# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

### v1

@@ 4. Automorphic forms and Fourier coefficients
This is an adelic reformulation of the classically well-known correspondence between the ideal classes and the ${\bf SL}_{2}({\mathbb{Z}})$-equivalence classes of integral binary quadratic forms. The proof is straight-forward. ∎
@@
(4.12) ${\bf R}(\varphi,E,\Lambda):=\sum_{[u]\in{\rm Cl}(E)}a_{\varphi}([T_{\theta}(u)])\,\Lambda(u)^{-1}.$
@@
Let $\varphi\in S_{\varrho}({\bf K}_{0}(N))$.
@@
    $B^{T_{\theta},\Lambda}(\varphi;g_{\infty})=w_{D}^{-1}\,{\bf B}_{\varrho}^{T_{\theta}}(g_{\infty})({\bf R}(\varphi,E,\Lambda)).$
@@
    We have
@@
    (4.14) $\varrho(\delta)(B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta}))$ $=B^{-T_{\theta}{\mathsf{s}}(\sigma),\Lambda}(\varphi;b_{\mathbb{R}}^{\theta}),$
    where $\sigma:=\left[\begin{smallmatrix}1&t\\ 0&-1\end{smallmatrix}\right]$.
@@
(i) The set ${\mathbb{A}}_{E}^{\times}/E^{\times}E_{\mathbb{R}}^{\times}$ is decomposed to a disjoint union of $\widehat{\mathfrak{o}}_{E}^{\times}$-orbits $[u]\widehat{\mathfrak{o}}_{E}^{\times}$ for different classes $[u]\in{\rm Cl}(E)$. Let $\mu$ denote the quotient measure on ${\mathbb{A}}_{E}^{\times}/E^{\times}E_{\mathbb{R}}^{\times}$. For $\tau\in{\mathbb{A}}_{E}^{\times}$, set $\mathsf{m}_{\theta}(\tau):=\mathsf{m}(I_{\theta}(\tau),{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau))$. Then by (3.6) and (4.6),
@@
$={\bf B}_{\varrho}^{T_{\theta}}(g_{\infty})\Bigl{(}\sum_{[u]\in{\rm Cl}(E)}\int_{[u]\widehat{\mathfrak{o}}_{E}^{\times}}a_{\varphi}(T_{\theta};\mathsf{m}_{\theta}(u))\,\Lambda(u)^{-1}\,{\rm{d}}\mu(u)\Bigr{)}.$
@@
$\sum_{[u]\in{\rm Cl}(E)}\mu([u]\widehat{\mathfrak{o}}_{E}^{\times})\,a_{\varphi}(T_{\theta};\mathsf{m}_{\theta}(u))\,\Lambda(u)^{-1}.$
@@
It remains to compute $\mu([u]\widehat{\mathfrak{o}}_{E}^{\times})$. Since any fiber of the natural surjection from $\widehat{\mathfrak{o}}_{E}^{\times}$ onto $[u]\widehat{\mathfrak{o}}_{E}^{\times}$ has a simply transitive action of the global unit group $\mathfrak{o}_{E}^{\times}$, which coincides with the set of roots of unity in $E^{\times}$, we have $\mu([u]\widehat{\mathfrak{o}}_{E}^{\times})=(1/\#\mathfrak{o}_{E}^{\times})=1/w_{D}$.
@@ 4.1. Sign condition
The Galois group ${\rm Gal}(E/{\mathbb{Q}})$ acts on ${\rm Cl}(E)$ naturally, hence on the orbit space ${{\mathcal{Q}}}_{\rm prim}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ through the bijection in Lemma 4.3(iii). The following lemma describes the conjugate action on ${{\mathcal{Q}}}_{\rm prim}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ explicitly. Recall the element $\sigma\in{\bf GL}_{2}({\mathbb{Z}})$ defined in Lemma 4.4 (ii).
@@
For $u\in{\rm Cl}_{E}$, the element $T_{\theta}(\bar{u})$ is ${\bf SL}_{2}({\mathbb{Z}})$-equivalent to $-T_{\theta}(u)\,{\mathsf{s}}\left(\delta\right)$. In other words, the conjugate action on ${\mathcal{Q}}_{\rm prim}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ is induced by the map
$T=\left[\begin{smallmatrix}a&2^{-1}b\\ 2^{-1}b&c\end{smallmatrix}\right]\longmapsto-T\,{\mathsf{s}}(\delta)=\left[\begin{smallmatrix}a&-2^{-1}b\\ -2^{-1}b&c\end{smallmatrix}\right].$
@@
By the decomposition of $I_{\theta}(u)$ for $u\in{\mathbb{A}}_{E}^{\times}$ in (4.11), we may take $\gamma_{\bar{u}}=\sigma\gamma_{u}\sigma^{-1}$. Thus,
$T_{\theta}(\bar{u})=T_{\theta}{\mathsf{s}}(\gamma_{\bar{u}})=T_{\theta}{\mathsf{s}}\left(\sigma\gamma_{u}\sigma^{-1}\right).$
A computation shows $T_{\theta}{\mathsf{s}}\left(\sigma\right)=-T_{\theta}$. Hence,
@@
For $\Lambda\in\widehat{{\rm Cl}(E)}$, recall its conjugate $\Lambda^{\dagger}\in\widehat{{\rm Cl}(E)}$ (see (2.25)).
@@
Let $\varphi\in S_{\varrho}({\bf K}_{0}(N))$ and $\Lambda\in\widehat{{\rm Cl}(E)}$ and suppose $B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta})\not=0$. Then,
@@ 4.2. Automorphic representations
$(\varphi|\varphi_{1})_{L^{2}}:=\int_{{\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})}\varphi(g)\overline{\varphi_{1}(g)}\,{\rm{d}}g,\qquad\varphi,\,\varphi_{1}\in V_{\pi}).$
@@ 4.3. Basic assumptions
From now on, we fix a pair $(\lambda=(l_{1},l_{2}),N)\in({\mathbb{Z}}^{2})_{\rm dom}\times{\mathbb{Z}}_{>0}$ and a character $\mu=\otimes_{p\leqslant\infty}\mu_{p}:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}{\mathbb{R}}^{\times}\to{\mathbb{C}}^{1}$ with $M:={\rm cond}(\mu)$ satisfying the following conditions:
@@
    All prime divisors of $NM$ are inert in $E/{\mathbb{Q}}$.
As usual, the character $\mu$ induces an even primitive Dirichlet character $\tilde{\mu}:({\mathbb{Z}}/M{\mathbb{Z}})^{\times}\rightarrow{\mathbb{C}}^{1}$ in such a way that $\mu(u)=\tilde{\mu}(a)^{-1}$ for all $u\in\widehat{\mathbb{Z}}^{\times}$ and $a\in{\mathbb{Z}}$ with $u-a\in M\widehat{\mathbb{Z}}$ ; thus $\mu_{p}(p)=\tilde{\mu}(p)$ for primes $p\nmid M$, and
@@
We remark that the conditions listed above (when $\mu={\bf 1}$ and $l_{1}=l_{2}$) are also imposed in the most part of [7]. Let $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm SK}$ denote the set of $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ that is the Saito-Kurokawa lift ([29]) of an cuspidal automorphic representation of ${\bf PGL}_{2}({\mathbb{A}})$, which is locally described by [38] (see also [39]). From [38, §4], the set $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm SK}=\varnothing$ unless $l_{1}=l_{2}$, i.e., $V_{\varrho}$ is one dimensional. We note that all the representations $\pi\cong\otimes_{p}\pi_{p}$ in $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)\setminus\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm SK}$ are non-CAP (cf. [42], [18, §3.5], [34, Corollary 4.5]); then, by [48], $\pi_{p}$ is tempered for all $p\nmid N$. Moreover, by invoking [3] (see also [40]) , any $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)\setminus\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm SK}$ is either a Yoshida lift, i.e., there exists a pair of irreducible cuspidal automorphic representations $(\sigma_{1},\sigma_{2})$ of ${\bf GL}_{2}({\mathbb{A}})$ such that $L(s,\pi)=L(s,\sigma_{1})L(s,\sigma_{2})$, or a “general type”, that is, there exists an irreducible cuspidal automorphic representation $\Pi$ of ${\bf GL}_{4}({\mathbb{A}})$ such that $L(s,\Pi,\wedge^{2})$ has a pole at $s=1$ and $L(s,\pi)=L(s,\Pi)$. Let $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm Y}$ (resp. $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm G}$) be the set of $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)$ that is a Yoshida lift (resp. a general type). Then,
@@
Let $N_{\pi}$ be the product of primes $p<\infty$ such that $\pi_{p}$ is not spherical. By condition (ii) in §4.2, we have $N_{\pi}\mid N$. Due to (A-ii), the representations $\pi_{p}$ for $p\mid N_{\pi}$ have non-zero ${\bf K}_{0}(p{\mathbb{Z}}_{p})$-fixed vectors, which, in turn, implies that $\pi_{p}$ is Iwahori spherical so that their isomorphism classes are listed in [37, Table 3] (For extended tables, we refer to [35, Appendix].) Note that in [37] a different symplectic form is used to define the symplectic group; up to the adjustment for this difference, the group $P_{1}$ in [37] is our ${\bf K}_{0}(p{\mathbb{Z}}_{p})$. In [35], all irreducible admissible representations of $\mathsf{G}({\mathbb{Q}}_{p})$ that admit local Bessel models are classified, and the result is conveniently summarized in [32, Table 2]. Here is a summary of what is available for our $\pi$ (as in §4.3):

### v2

@@ 4. Automorphic forms and Fourier coefficients
Since (i) and (ii) are proved directly, we only give some detail of (iii). For $u,u^{\prime}\in{\mathbb{A}}_{E}^{\times}$, let $I_{\theta}(u)=\gamma_{u}h_{u}\kappa_{u}$ and $I_{\theta}(u^{\prime})=\gamma_{u^{\prime}}h_{u^{\prime}}\kappa_{u^{\prime}}$ be as in (4.11). If $T_{\theta}(u)$ and $T_{\theta}(u^{\prime})$ are ${\bf\Gamma}_{0}({\mathcal{M}})$-equivalent, by Lemma 2.1, there exist $\delta_{0}\in{\bm{\Gamma}}_{0}({\mathcal{M}})$ and $r\in E^{\times}$ such tha $\gamma_{u^{\prime}}=I_{\theta}(r)\gamma_{u}\delta_{0}$. From this and noting that $T_{\theta}=T_{\theta}{\mathsf{s}}(I_{\theta}(u))=T_{\theta}{\mathsf{s}}(I_{\theta}(u^{\prime}))$ and $I_{\theta}^{-1}({\bf\Gamma}_{0}({\mathcal{M}}\widehat{\mathbb{Z}}))\subset U({\mathcal{M}})$, we find that there exist $c_{\rm{\bf{f}}}\in U({\mathcal{M}})$, $c_{\infty}\in E_{\infty}^{\times}$ such that $\kappa_{u^{\prime}}=\delta_{0,{\rm{\bf{f}}}}^{-1}\kappa_{u}I_{\theta}(c_{\rm{\bf{f}}})$, $h_{u^{\prime}}=\delta_{0,\infty}^{-1}h_{u}I_{\theta}(c_{\infty})$, which yields
$I_{\theta}(u^{\prime})=I_{\theta}(r)\gamma_{u}h_{u}\kappa_{u}I_{\theta}(c_{\infty}c_{\rm{\bf{f}}})=I_{\theta}(urc_{\rm{\bf{f}}}c_{\infty}).$
Since $I_{\theta}$ is injective, we conclude that $u^{\prime}$ $u$ are ${\mathbb{A}}^{\times}E^{\times}E_{\infty}^{\times}U({\mathcal{M}})$-equivalent, completing the proof of injectivity of the map. To show the surjectivity, take an element $T\in{\mathcal{Q}}_{{\mathcal{M}}}^{+}(D)$. By Witt’s theorem and Lemma 2.1, $T=T_{\theta}{\mathsf{s}}(\gamma)$ for some $\gamma\in{\bf GL}_{2}({\mathbb{Q}})$. On the other hand, for each place $p$, we can prove that there exists an element $g_{p}\in{\bm{{\bf GL}}}_{2}({\mathbb{Z}}_{p})$ (resp. $g_{p}\in{\bm{\Gamma}}_{0}({\mathcal{M}}{\mathbb{Z}}_{p})$) for $p\nmid{\mathcal{M}}$ (resp. $p\mid{\mathcal{M}}$) such that $T=T_{\theta}{\mathsf{s}}(g_{p})$ in ${\bf GL}_{2}({\mathbb{Q}}_{p})$. (Indeed, the case $p\mid{\mathcal{M}}$ is due to a straightforward calculation. When $p\nmid{\mathcal{M}}$, we use [44, Lemma 2-4], noting that $T,T_{\theta}\in({\mathcal{Q}}_{{\mathbb{Z}}_{p}})_{\rm prim}$, to find $u_{p}\in E_{p}^{\times}$ and $g_{p}\in{\bf GL}_{2}({\mathbb{Z}}_{p})$ such that $\gamma=I_{\theta}(u_{p})g_{p}$. Then, $g_{p}$ works.) Moreover, the condition $\det g_{\infty}>0$ follows from the positive definiteness of $T$ and $T_{\theta}$. Therefore, by Lemma 2.1, there exists an element $u\in{\mathbb{A}}_{E}^{\times}$ such that
$I_{\theta}(u)=\gamma h\kappa,\quad(h:=g_{\infty}^{-1}\in{\bf GL}_{2}({\mathbb{R}})^{\circ},\kappa:=(g_{p}^{-1})_{p<\infty}\in{\bf\Gamma}_{0}({\mathcal{M}}\widehat{\mathbb{Z}})).$
This shows that $T_{\theta}(u)=T_{\theta}{\mathsf{s}}(\gamma)=T$. ∎
@@
(4.12) ${\bf R}(\varphi,E,\Lambda):=\sum_{[u]\in{\rm Cl}_{{\mathcal{M}}}^{0}(E)}a_{\varphi}([T_{\theta}(u)])\,\Lambda(u)^{-1}.$
@@
Let $\varphi\in S_{\varrho}({\bf K}_{0}(N))$ and $\Lambda\in\widehat{{\rm Cl}_{{\mathcal{M}}}^{0}(E)}_{{\rm prim}}$.
@@
    $B^{T_{\theta},\Lambda}(\varphi;g_{\infty})=w_{D,{\mathcal{M}}}^{-1}\,{\bf B}_{\varrho}^{T_{\theta}}(g_{\infty})({\bf R}(\varphi,E,\Lambda))$
    where $w_{D,N}:=\#(U({\mathcal{M}})\cap{\mathfrak{o}}_{E}^{\times})$.
@@
    Set $\sigma:=\left[\begin{smallmatrix}1&t\\ 0&-1\end{smallmatrix}\right]$. Then,
@@
    (4.14) $\varrho(\delta)(B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta}))$ $=B^{-T_{\theta}{\mathsf{s}}(\sigma),\Lambda}(\varphi;b_{\mathbb{R}}^{\theta}).$
@@
(i) The set ${\mathbb{A}}_{E}^{\times}/E^{\times}E_{\mathbb{R}}^{\times}$ is decomposed to a disjoint union of ${\mathbb{A}}_{{\rm{\bf{f}}}}^{\times}U({\mathcal{M}})$-orbits $[u]{\mathbb{A}}_{{\rm{\bf{f}}}}^{\times}U({\mathcal{M}})$ for different classes $[u]\in{\rm Cl}_{{\mathcal{M}}}^{0}(E)$. Let $\mu$ denote the quotient measure on ${\mathbb{A}}_{E}^{\times}/E^{\times}E_{\mathbb{R}}^{\times}$. For $\tau\in{\mathbb{A}}_{E}^{\times}$, set $\mathsf{m}_{\theta}(\tau):=\mathsf{m}(I_{\theta}(\tau),{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau))$. Then by (3.7) and (4.6),
@@
$={\bf B}_{\varrho}^{T_{\theta}}(g_{\infty})\Bigl{(}\sum_{[u]\in{\rm Cl}_{{\mathcal{M}}}^{0}(E)}\int_{[u]{\mathbb{A}}_{{\rm{\bf{f}}}}^{\times}U({\mathcal{M}})}a_{\varphi}(T_{\theta};\mathsf{m}_{\theta}(u))\,\Lambda(u)^{-1}\,{\rm{d}}\mu(u)\Bigr{)}.$
@@
$\sum_{[u]\in{\rm Cl}_{{\mathcal{M}}}^{0}(E)}\mu([u]{\mathbb{A}}_{{\rm{\bf{f}}}}^{\times}U({\mathcal{M}}))\,a_{\varphi}(T_{\theta};\mathsf{m}_{\theta}(u))\,\Lambda(u)^{-1}.$
@@
It remains to compute $\mu([u]{\mathbb{A}}_{{\rm{\bf{f}}}}^{\times}U({\mathcal{M}}))$. Since any fiber of the natural surjection from ${\mathbb{A}}_{{\rm{\bf{f}}}}^{\times}U({\mathcal{M}})$ onto $[u]{\mathbb{A}}_{{\rm{\bf{f}}}}^{\times}U({\mathcal{M}})$ has a simply transitive action of the global unit group $\mathfrak{o}_{E}^{\times}$, we have $\mu([u]{\mathbb{A}}_{{\rm{\bf{f}}}}^{\times}U({\mathcal{M}}))=1/\#(U({\mathcal{M}})\cap{\mathfrak{o}}_{E}^{\times})=1/w_{D,N}$.
@@ 4.1. Sign condition
The Galois group ${\rm Gal}(E/{\mathbb{Q}})$ acts on ${\rm Cl}_{{\mathcal{M}}}^{0}(E)$ naturally, hence on the orbit space ${{\mathcal{Q}}}_{{\mathcal{M}}}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ through the bijection in Lemma 4.3(iii). The following lemma describes the conjugate action on ${{\mathcal{Q}}}_{{\mathcal{M}}}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ explicitly. Recall the element $\sigma\in{\bf GL}_{2}({\mathbb{Z}})$ defined in Lemma 4.4 (ii).
@@
For $u\in{\rm Cl}_{\mathcal{M}}^{0}(E)$, the element $T_{\theta}(\bar{u})$ is ${\bf\Gamma}_{0}({\mathcal{M}})$-equivalent to $-T_{\theta}(u)\,{\mathsf{s}}\left(\delta\right)$. In other words, the conjugate action on ${\mathcal{Q}}_{{\mathcal{M}}}^{+}(D)/{\bf\Gamma}_{0}({\mathcal{M}})$ is induced by the map $T\mapsto-T\,{\mathsf{s}}(\delta)$.
@@
By the decomposition of $I_{\theta}(u)$ for $u\in{\mathbb{A}}_{E}^{\times}$ in (4.11), we may take $\gamma_{\bar{u}}=\sigma\gamma_{u}\sigma^{-1}$. Thus, $T_{\theta}(\bar{u})=T_{\theta}{\mathsf{s}}(\gamma_{\bar{u}})=T_{\theta}{\mathsf{s}}\left(\sigma\gamma_{u}\sigma^{-1}\right)$. A computation shows $T_{\theta}{\mathsf{s}}\left(\sigma\right)=-T_{\theta}$. Hence,
@@
For $\Lambda\in\widehat{{\rm Cl}_{{\mathcal{M}}}^{0}(E)}_{\rm prim}$, recall its conjugate $\Lambda^{\dagger}\in\widehat{{\rm Cl}_{{\mathcal{M}}}^{0}(E)}_{\rm prim}$ (see (2.24)).
@@
Let $\varphi\in S_{\varrho}({\bf K}_{0}(N))$ and $\Lambda\in\widehat{{\rm Cl}_{{\mathcal{M}}}^{0}(E)}_{\rm prim}$ and suppose $B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta})\not=0$. Then,
@@ 4.2. Automorphic representations
$(\varphi\mid\varphi_{1})_{L^{2}}:=\int_{{\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})}\varphi(g)\overline{\varphi_{1}(g)}\,{\rm{d}}g,\qquad\varphi,\,\varphi_{1}\in V_{\pi}.$
@@ 4.3. Basic assumptions
From now on, we fix a pair $(\lambda=(l_{1},l_{2}),N,M,{\mathcal{M}})\in({\mathbb{Z}}^{2})_{\rm dom}\times{\mathbb{Z}}_{>0}\times{\mathbb{Z}}_{>0}\times{\mathbb{Z}}_{>0}$ satisfying the following conditions:
@@
    All prime divisors of $NM{\mathcal{M}}$ are inert in $E/{\mathbb{Q}}$.
We fix characters $\Lambda\in\widehat{{\rm Cl}_{{\mathcal{M}}}^{0}(E)}_{{\rm prim}}$ and $\mu=\otimes_{p\leqslant\infty}\mu_{p}\in\widehat{{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}{\mathbb{R}}_{>0}}$ with $M:={\rm cond}(\mu)$ as well. As usual, the character $\mu$ induces a primitive Dirichlet character $\tilde{\mu}:({\mathbb{Z}}/M{\mathbb{Z}})^{\times}\rightarrow{\mathbb{C}}^{1}$ in such a way that $\mu(u)=\tilde{\mu}(a)^{-1}$ for all $u\in\widehat{\mathbb{Z}}^{\times}$ and $a\in{\mathbb{Z}}$ with $u-a\in M\widehat{\mathbb{Z}}$ ; thus $\mu_{p}(p)=\tilde{\mu}(p)$ for primes $p\nmid M$, and
@@
We remark that the conditions listed above (when $\mu={\bf 1}$ and $l_{1}=l_{2}$) are also imposed in the most part of [7]. Let $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm SK}$ denote the set of $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ that is the Saito-Kurokawa lift ([30]) of a cuspidal automorphic representation of ${\bf PGL}_{2}({\mathbb{A}})$, which is locally described by [39] (see also [40]). From [39, §4], the set $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm SK}=\varnothing$ unless $l_{1}=l_{2}$, i.e., $V_{\varrho}$ is one dimensional. We note that all the representations $\pi\cong\otimes_{p}\pi_{p}$ in $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)\setminus\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm SK}$ are non-CAP (cf. [43], [17, §3.5], [35, Corollary 4.5]); then, by [49], $\pi_{p}$ is tempered for all $p\nmid N$. Moreover, by invoking [3] (see also [41]) , any $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)\setminus\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm SK}$ is either a Yoshida lift, i.e., there exists a pair of irreducible cuspidal automorphic representations $(\sigma_{1},\sigma_{2})$ of ${\bf GL}_{2}({\mathbb{A}})$ such that $L(s,\pi)=L(s,\sigma_{1})L(s,\sigma_{2})$, or a “general type”, that is, there exists an irreducible cuspidal automorphic representation $\Pi$ of ${\bf GL}_{4}({\mathbb{A}})$ such that $L(s,\Pi,\wedge^{2})$ has a pole at $s=1$ and $L(s,\pi)=L(s,\Pi)$. Let $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm Y}$ (resp. $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm G}$) be the set of $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)$ that is a Yoshida lift (resp. a general type). Then,
@@
Let $N_{\pi}$ be the product of primes $p<\infty$ such that $\pi_{p}$ is not spherical. By condition (ii) in §4.2, we have $N_{\pi}\mid N$. Due to (A-ii), the representations $\pi_{p}$ for $p\mid N_{\pi}$ have non-zero ${\bf K}_{0}(p{\mathbb{Z}}_{p})$-fixed vectors, which, in turn, implies that $\pi_{p}$ is Iwahori spherical so that their isomorphism classes are listed in [38, Table 3] (For extended tables, we refer to [36, Appendix].) Note that in [38] a different symplectic form is used to define the symplectic group; up to the adjustment for this difference, the group $P_{1}$ in [38] is our ${\bf K}_{0}(p{\mathbb{Z}}_{p})$. In [36], all irreducible admissible representations of $\mathsf{G}({\mathbb{Q}}_{p})$ that admit local Bessel models are classified, and the result is conveniently summarized in [33, Table 2]. Here is a summary of what is available for our $\pi$ (as in §4.3):